### v4

@@ 4.2 Avoidance of strictly saddle points
Proof.
@@
implying
$|\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})-\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})|\leq|\hat{z}_{i}-\tilde{z}_{i}|+|\hat{w}_{i}-\tilde{w}_{i}|.$
(c) If $|\hat{z}_{i}|>\hat{w}_{i}$ and $|\tilde{z}_{i}|>\tilde{w}_{i}$, then
$|\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})-\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})|=|\text{sign}(\hat{z}_{i})(|\hat{z}_{i}|-\hat{w}_{i})-\text{sign}(\tilde{z}_{i})(|\tilde{z}_{i}|-\tilde{w}_{i})|.$
If $\text{sign}(\hat{z}_{i})=\text{sign}(\tilde{z}_{i})$, $|\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})-\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})|=|(|\hat{z}_{i}|-\hat{w}_{i})-(|\tilde{z}_{i}|-\tilde{w}_{i})|\leq|\hat{z}_{i}-\tilde{z}_{i}|+|\hat{w}_{i}-\tilde{w}_{i}|$. If $\text{sign}(\hat{z}_{i})=-\text{sign}(\tilde{z}_{i})$, $|\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})-\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})|=(|\hat{z}_{i}|-\hat{w}_{i})+(|\tilde{z}_{i}|-\tilde{w}_{i})\leq|\hat{z}_{i}|+|\tilde{z}_{i}|=|\hat{z}_{i}-\tilde{z}_{i}|.$
In all three cases, we show
$|\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})-\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})|\leq|\hat{z}_{i}-\tilde{z}_{i}|+|\hat{w}_{i}-\tilde{w}_{i}|.$
Therefore, by (A.1),
$\|\mathbb{S}_{\hat{w}}(\hat{w})-\mathbb{S}_{\tilde{w}}(\tilde{w})\|\leq\|\hat{z}-\tilde{z}\|+\|\hat{w}-\tilde{w}\|.$

### v5

@@ Appendix A Proof supplementary
Proof A.2.
We consider the following three cases.
@@
implying $|\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})-\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})|\leq|\hat{z}_{i}-\tilde{z}_{i}|+|\hat{w}_{i}-\tilde{w}_{i}|.$
(c) If $|\hat{z}_{i}|>\hat{w}_{i}$ and $|\tilde{z}_{i}|>\tilde{w}_{i}$, then $|\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})-\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})|=|\text{sign}(\hat{z}_{i})(|\hat{z}_{i}|-\hat{w}_{i})-\text{sign}(\tilde{z}_{i})(|\tilde{z}_{i}|-\tilde{w}_{i})|.$ If $\text{sign}(\hat{z}_{i})=\text{sign}(\tilde{z}_{i})$, $|\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})-\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})|=|(|\hat{z}_{i}|-\hat{w}_{i})-(|\tilde{z}_{i}|-\tilde{w}_{i})|\leq|\hat{z}_{i}-\tilde{z}_{i}|+|\hat{w}_{i}-\tilde{w}_{i}|$. If $\text{sign}(\hat{z}_{i})=-\text{sign}(\tilde{z}_{i})$, $|\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})-\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})|=(|\hat{z}_{i}|-\hat{w}_{i})+(|\tilde{z}_{i}|-\tilde{w}_{i})\leq|\hat{z}_{i}|+|\tilde{z}_{i}|=|\hat{z}_{i}-\tilde{z}_{i}|.$
In all the three cases, we show $|\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})-\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})|\leq|\hat{z}_{i}-\tilde{z}_{i}|+|\hat{w}_{i}-\tilde{w}_{i}|.$ Therefore, $\|\mathbb{S}_{\hat{w}}(\hat{w})-\mathbb{S}_{\tilde{w}}(\tilde{w})\|\leq\|\hat{z}-\tilde{z}\|+\|\hat{w}-\tilde{w}\|.$